solve the canchy-Euler inital volue Problem. x²y )) -sxy +5420, Y(1)=1, Y(1)=9 Show all work
Answer #1

pomy Y CHE HA 45324 rattrups suhtes Ginen BJ Ded y c12=1 OS! =A+B YLD I 1 A+B Q hoold Cinen gen=9 y co= a +5B V A+5B AB=1 18